A chemist is mixing a solution containing 15% bromine with a second solution containing 28% bromine. She wishes to prepare 60 oz. of a solution containing 21% bromine. How much of each of the solutions should she use? [Round to the nearest tenth of an ounce.]
Answer by stanbon(75887)

A chemist is mixing a solution containing 15% bromine with a second solution containing 28% bromine. She wishes to prepare 60 oz. of a solution containing 21% bromine. How much of each of the solutions should she use?
---------------------------
Equation:
0.15x + 0.28(60-x) = 0.21*60
----
Multiply thru by 100:
15x + 28*60 - 28x = 21*60
-13x = -7*60
x = (7/13)60
x = 32.3 oz (amt. of 15% solution needed)
------
60-x = 27.7 oz (amt. of 28% solution needed)
=====================
Cheers,
Stan H.